OpenStudy (inowalst):

If \(f(x)=x^2+\sqrt[3]{x}\), what is the value of \(f(-8)\) ?

OpenStudy (inowalst):

I dont get this. @tkhunny

OpenStudy (tkhunny):

Have you considered entering -8 for every occurrence of 'x' in the formulation? That IS what the notation means. What is the CUBE root of -8? Not the square root.

OpenStudy (inowalst):

-2

OpenStudy (inowalst):

\(f(x)=-8^2-2\)

OpenStudy (tkhunny):

\((-8)^{2} + \sqrt[3]{-8} = 64 - 2\) You have it.

OpenStudy (inowalst):

Oh okay. Thank you.

OpenStudy (tkhunny):

Careful with the notation \(-8^{2}\). It may or may not mean what you want. With the parentheses, it is unambiguous.

OpenStudy (inowalst):

It's a math problem from my SAT book and that's how it says it..

OpenStudy (tkhunny):

It says \(x^{2}\). This implies \((-8)^{2}\). If you write \(-8^{2}\), it MAY mean the same thing. Computer languages have conventions for this sort of thing - called "precedence".

OpenStudy (inowalst):

Ohhh. Then, yes. I meant the -8 to be in parenthesis.